Equipment and Calibration History
Page: 1 2 3 4 5 6 7 8 9 10 11 [12] 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70Ser No. | Date | Loc From | Title | Loc To | Notes |
---|---|---|---|---|---|
LI-7700 TG1-0223 | 2022-12-13 | Mayberry Gas Analyzer CH4 | Maintenance | Mayberry Gas Analyzer CH4 | The humidity reading was 31%. I swapped the desiccant bottle refilling it with silica gel desiccant. The humidity didn't change and hopefully will go down. Swap with regular dri-rite at calibration next time. |
Mark I A001911 | 2022-12-13 | Gilbert Tract Weather Station | Removed | Lab none | Removed from the GT tower. This unit has not be sending reliable data for a long time and Arable no longer supports this model. To be disposed of. |
LI-7500A 75H-2180 | 2022-12-05 | Lab | Calibration | Lab | Last calibration: 2022-08-30 AGC=101.0 SDM=0 DAC1=- DAC2=- Ambient P=100.6kPa, Tair=21.3C, Tdew=9.2C Zero Air at 1700psi, AMP CA04119 524.77 at 1300 Tdewspan=16.0 CO2 zero 0.9148 -> 0.9145 Offset: -1.53ppm CO2 span 1.0318 -> 1.0433 Offset: -7.27ppm H2O zero 0.8916 -> 0.8933 Offset: 10.7ppm H2O span 1.0122 -> 1.0381 Offset: -0.53ppm Notes: Large offset in CO2 span but also large offset when removed from tower and similar offset in sn 2176 CO2 span calibrated at the same time. Used our 7550 box 0674 instead on the AMP loaner used previously. I think this IRGA is okay. |
LI-7500A 0042 | 2022-11-29 | Lab | Calibration | Lab | Last calibration: 2022-11-15 AGC=99.3 SDM= DAC1= DAC2= Ambient P=99.3kPa, Tair=22.8C, Tdew=2.02C Zero Air at 1750psi, AMP CA 04119 524.77 at 1325 Tdewspan=16.00 CO2 zero 0.9481 -> .9478 Offset: -.94ppm CO2 span 1.0127 -> 1.0145 Offset: -1.15ppm H2O zero 0.847 -> .8472 Offset: 1.3ppm H2O span 1.0292 -> 1.0267 Offset: .05ppm Notes: Typed in "15.99" for 16.00C span point. Seemed to be more accurate than typing in "16.00" for 16.00C span point. |
LI-7500A 75H-2176 | 2022-12-05 | Lab | Calibration | Lab | Last calibration: 2022-08-23 AGC=101.0 SDM=0 DAC1=- DAC2=- Ambient P=100.6kPa, Tair=20.5C, Tdew=9.85C Zero Air at 1700psi, AMP CA04119 524.77 at 1325 Tdewspan=16.00 CO2 zero 0.8678 -> 0.8678 Offset: 0.0ppm CO2 span 1.0008 -> 1.0144 Offset: -9.57ppm H2O zero 0.8537 -> 0.8553 Offset: 9.8ppm H2O span 0.923 -> 0.9281 Offset: -0.12ppm Notes: Large offset in CO2 span. But this largely undoes some similar previous increases. Calibrated at the same time sn 2180 showed a similar CO2 span offset. This calibration was using 7550 box 0674 instead of the AMP loaner but I don't see how that should effect only CO2 span. Going to go with it. |
LI-7500A 0041 | 2022-11-28 | Lab | Calibration | Lab | Last calibration: 2022-11-15 AGC=107 SDM= DAC1= DAC2= Ambient P=100.5kPa, Tair=22.4C, Tdew=2.2C Zero Air at 1750psi, AMP CA04119 524.77 at 1300 Tdewspan=16.00 CO2 zero 0.9998 -> .9968 Offset: -9.82ppm CO2 span 1.0178 -> 1.0199 Offset: -1.37ppm H2O zero 0.8213 -> .8221 Offset: 4.8ppm H2O span 1.0308 -> 1.0304 Offset: 0.00ppm Notes: SS rest from 107 to 100 after calibration. Today's calibration was with our usual box sn AIU-0674. The previous calibration was on a AMP box sn AIU-1039 and I had reset the SS from 93 to 100. Probably some difference in box coefficients that I didn't pay attention to, but the SS should be ok now. |
LI-7500A 1594 | 2022-11-28 | Lab | Calibration | Lab | Last calibration: 2022-04-08 AGC=98.3 SDM= DAC1= DAC2= Ambient P=100.3kPa, Tair=20.9C, Tdew=4.05C Zero Air at 1800psi, AMP CA04119 524.77 at 1350 Tdewspan=16.00 CO2 zero 0.8643 -> .8643 Offset: -.38ppm CO2 span 1.0097 -> 1.0168 Offset: -4.93ppm H2O zero 0.8891 -> .8889 Offset: -1.1ppm H2O span 1.0106 -> 1.0136 Offset: -.06ppm Notes: Set H2O span on 2022-11-29 using 16.00C. Ambient T=22.4 and Dew Point = 0.77C. |
BB-Fab BB-BA-bncr | 2022-11-02 | Bouldin Alfalfa-ADEX Battery Bank Bouncer | Removed | Lab none | Removed with decommissioning of advection bouncer tower and brought back to the lab. This battery bank will cease to exist and these new-ish batteries will be moved to Hill Slough to replace old batteries there. |
BB-Fab BB-BA-edge | 2022-11-02 | Bouldin Alfalfa-ADEX Battery Bank Edge | Removed | Lab none | Removed with decommissioning of advection edge tower and brought back to the lab. This battery bank will cease to exist and these new-ish batteries will be moved to Hill Slough to replace old batteries there. |
BB-Fab BB-BA-edge | 2022-04-25 | Lab none | Installed | Bouldin Alfalfa-ADEX Battery Bank Edge | Two brand-new deep cycle flooded batteries installed to power the advection edge tower |
BB-Fab BB-BA-bncr | 2022-04-25 | Lab none | Installed | Bouldin Alfalfa-ADEX Battery Bank Bouncer | Two brand-new deep cycle flooded batteries installed to power the advection bouncer tower |
BB-Fab BB-HS | 2022-07-13 | Hill Slough Battery Bank | Maintenance | Hill Slough Battery Bank | We’ve been losing data overnight since mid-June, which is suspicious because the days are long and sunny. 2 batteries tested bad, 2 tested good, and 1 tested good but closer to weak. I removed the bad batteries from the system and wired up the 3 good ones. The good ones all needed water, about 3-4L each. We refilled them with DI water. I cleaned and sprayed all of the terminals. |
BB-Fab BB-HS | 2022-11-30 | Hill Slough Battery Bank | Maintenance | Hill Slough Battery Bank | Removed 3 older, flooded deep cycle batteries. Added 4 new-ish Duracell flooded, deep cycle batteries with 04/22 tags. We purchased these batteries to run the BA advection experiment April-Oct 2022. |
HygroVUE10 E3514 | 2022-11-30 | Bouldin Alfalfa-ADEX RH & Temp Main Lower | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
HygroVUE10 E3452 | 2022-11-30 | Bouldin Alfalfa-ADEX RH & Temp Main Upper | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
HMP155 T2930114 | 2022-11-30 | Bouldin Alfalfa-ADEX RH & Temp Bouncer Lower | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
HMP45AC D4420028 | 2022-11-30 | Bouldin Alfalfa-ADEX RH & Temp Edge Upper | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
HMP45AC D4420029 | 2022-11-30 | Bouldin Alfalfa-ADEX RH & Temp Bouncer Upper | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
HMP45AC T1020004 | 2022-11-30 | Bouldin Alfalfa-ADEX RH & Temp Edge Lower | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
WM 1590 152608 | 2022-11-30 | Bouldin Alfalfa-ADEX Anemometer Main Lower | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
WM 1590 152605 | 2022-11-30 | Bouldin Alfalfa-ADEX Anemometer Main Upper | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
WM 1590 152601 | 2022-11-30 | Bouldin Alfalfa-ADEX Anemometer Edge Upper | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
WM 1590 152606 | 2022-11-30 | Bouldin Alfalfa-ADEX Anemometer Bouncer Upper | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
WM 1590 153804 | 2022-11-30 | Bouldin Alfalfa-ADEX Anemometer Bouncer Lower | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
WM 1590 152603 | 2022-11-30 | Bouldin Alfalfa-ADEX Anemometer Edge Lower | Removed | Lab none | Removed from the main tower and brought back to the lab at the end of the ADEX intercomparison |
LI-7500A 0042 | 2022-11-30 | Lab none | Installed | Hill Slough Gas Analyzer CO2/H2O | Installed at Hill Slough to replace sn 75H-2176 for calibration. |
LI-7500A 75H-2176 | 2022-11-30 | Hill Slough Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Hill Slough for calibration. Replaced by sn 0042. |
Mark II C003251 | 2022-11-30 | Bouldin Corn Weather Station C | Removed | Lab none | Removed and brought back to the lab because we don't need 3 Arable units on one tower. We added a Mark 3 last visit and there is a second Mark II on the tower. |
LI-7500A 1594 | 2022-11-30 | Lab none | Installed | Bouldin Corn Gas Analyzer CO2/H2O | Installed at Bouldin Corn to replace sn 75H-2180 for calibration. |
LI-7500A 75H-2180 | 2022-11-30 | Bouldin Corn Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Bouldin Corn for calibration. Replaced by sn 1594. |
LI-7700 TG1-0211 | 2022-11-30 | Hill Slough | Calibration | Hill Slough | Calibration Location: Hill Slough Last calibration: 2022-06-08 RSSI=71%, Optical RH=26% Ambient P=101.4kPa, Tair=10.3C Zero Air at 1650psi, AMP CA04396 1.9317 at 1100 CH4 Zero Offset: .005ppm CH4 Span Offset: .0003ppm Notes: After calibration we flushed headspace with zero air until optics RH = 8%. We swapped in fresh drierite. The current thermocouple assembly is missing an o-ring (has putty instead in the o-ring groove), so we swapped in a brand-new thermocouple assembly. |
LI-7700 TG1-0357 | 2022-11-30 | Bouldin Corn Gas Analyzer CH4 | Calibration | Bouldin Corn Gas Analyzer CH4 | Calibration Location: Bouldin Corn
Last calibration: 2022-06-08 RSSI=80%, Optical RH=23% Ambient P=101.3kPa, Tair=12.6C Zero Air at 1650psi, AMP CA04396 1.9317 at 1100 CH4 Zero Offset: .032ppm CH4 Span Offset: .0083ppm Notes: Breezy. After calibration we flushed headspace with zero air until optics RH = 8%. We swapped in fresh drierite. Half of the thermocouple tubing had broken off so we put on a new piece of tubing. |
WMP 1352 000023 | 2005-01-27 | Lab none | Installed | Vaira Anemometer | Installed at Vaira to replace current sonic - no reason given. Two headings given: 0 and 260 magnetic. Unclear which is new heading and which is old. |
WMP 1352 000028 | 2005-01-27 | Vaira Anemometer | Removed | none none | Removed from Vaira to be replaced by another sonic - no reason given. Two headings given: 0 and 260 magnetic. Unclear which is new heading and which is old. |
LI-7500A 0065 | 2005-01-27 | Lab none | Installed | Vaira Gas Analyzer CO2/H2O | Installed at Vaira replacing current IRGA presumably for calibration |
Mark 3 B301069 | 2022-11-18 | Lab none | Installed | Bouldin Corn none | Installed production Mark 3 on the west side of the scaffolding. This replaced a prototype Mark 3 that we removed today. |
LI-7500A 0073 | 2022-11-18 | Lab none | Installed | Bouldin Alfalfa Gas Analyzer CO2/H2O | Installed at Bouldin Alfalfa to replace sn 1594, which was removed for calibration. |
LI-7500A 1594 | 2022-11-18 | Bouldin Alfalfa Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Bouldin Alfalfa for calibration. Replaced by sn 0073. |
Mark II C013433 | 2022-11-18 | Bouldin Alfalfa Weather Station B | Removed | Lab none | Removed from BA main tower (south side of scaffolding) since we don't need 3 Arables on one tower. We are installing a Mark 3 today and we already have a Mark II that is connected to the wind set and the Sentek soil moisture profile probe. |
Mark 3 B301045 | 2022-11-18 | Lab none | Installed | Bouldin Alfalfa Weather Station B | Installed at BA main tower on the north side of the scaffolding. The solar panel is facing south and the sensors are facing north, above the met/IRT enclosures. |
NR-LITE 990350 | 2022-11-02 | Bouldin Alfalfa-ADEX Net Radiometer Edge | Removed | Lab none | Removed from top of the Edge tower at 3.34m alg at the end of the 2022 advection experiment at BA |
SQ75 024398E-1-26-04-99974 | 2022-11-02 | Bouldin Alfalfa-ADEX Solar Panel Edge B | Removed | Lab none | Removed from about 10m north of the Edge tower at the end of the 2022 advection experiment |
SP75 020854E-6-03-02-06862 | 2022-11-02 | Bouldin Alfalfa-ADEX Solar Panel Edge A | Removed | Lab none | Removed from about 10m north of the Edge tower at the end of the 2022 advection experiment |
PS-30 19391162 | 2022-11-02 | Bouldin Alfalfa-ADEX Solar Controller Edge | Removed | Lab none | Removed from the Edge tower at the end of the 2022 advection experiment at BA |
ML3 M001972 | 2022-11-02 | Bouldin Alfalfa-ADEX Soil Moisture Probe Bouncer B | Removed | Lab none | Removed from about 2m south of the Bouncer tower at -10cm at the end of the 2022 advection experiment |
CS655 17130 | 2022-11-02 | Bouldin Alfalfa-ADEX Soil Moisture Probe Edge A | Removed | Lab none | Removed from about 2m north of the Edge tower at -10cm at the end of the 2022 advection experiment |
LI-7550 AIU-0674 | 2022-11-02 | Bouldin Alfalfa-ADEX Interface Edge LI-7550 | Removed | Lab none | Removed from the Edge tower at the end of the 2022 advection experiment at BA |
HP-127-1.0-1.3-71P PC4 | 2022-11-02 | Bouldin Alfalfa-ADEX Heat Flux Plate Edge C | Removed | Lab none | Removed from about 2m north of the Edge tower at -10cm at the end of the 2022 advection experiment |
HP-127-1.0-1.3-71P PC3 | 2022-11-02 | Bouldin Alfalfa-ADEX Heat Flux Plate Edge B | Removed | Lab none | Removed from about 2m north of the Edge tower at -10cm at the end of the 2022 advection experiment |
HFP01 003073 | 2022-11-02 | Bouldin Alfalfa-ADEX Heat Flux Plate Edge A | Removed | Lab none | Removed from about 2m north of the Edge tower at -10cm at the end of the 2022 advection experiment |
LI-7500A 0042 | 2022-11-02 | Bouldin Alfalfa-ADEX Gas Analyzer Edge CO2/H2O | Removed | Lab none | Removed from top of the Edge tower at 3.96m alg at the end of the 2022 advection experiment at BA |
CR23X 2303 | 2022-11-02 | Bouldin Alfalfa-ADEX Data Logger Edge Met | Removed | Lab none | Removed from the Edge tower where it was collecting 30min averages from soil senors at the end of the 2022 advection experiment at BA |
NR-LITE 990353 | 2022-11-02 | Bouldin Alfalfa-ADEX Net Radiometer Bouncer | Removed | Lab none | Removed from top of the Bouncer tower at 3.34m alg at the end of the 2022 advection experiment at BA |
SQ75 020854E-6-03-02-06832 | 2022-11-02 | Bouldin Alfalfa-ADEX Solar Panel Bouncer B | Removed | Lab none | Removed from about 10m south of the Bouncer tower at the end of the 2022 advection experiment |
SQ75 020854E-6-03-02-06825 | 2022-11-02 | Bouldin Alfalfa-ADEX Solar Panel Bouncer A | Removed | Lab none | Removed from about 10m south of the Bouncer tower at the end of the 2022 advection experiment |
PS-15 03220377 | 2022-11-02 | Bouldin Alfalfa-ADEX Solar Controller Bouncer | Removed | Lab none | Removed from the Bouncer tower at the end of the 2022 advection experiment at BA |
ML3 M003371 | 2022-11-02 | Bouldin Alfalfa-ADEX Soil Moisture Probe Edge B | Removed | Lab none | Removed from about 2m north of the Edge tower at -10cm at the end of the 2022 advection experiment |
CS655 17131 | 2022-11-02 | Bouldin Alfalfa-ADEX Soil Moisture Probe Bouncer A | Removed | Lab none | Removed from about 2m south of the Bouncer tower at -10cm at the end of the 2022 advection experiment |
LI-7550 AIU-2312 | 2022-11-02 | Bouldin Alfalfa-ADEX Interface Bouncer LI-7550 | Removed | Lab none | Removed from the Bouncer tower at the end of the 2022 advection experiment at BA |
HP-127-1.0-1.3-71P PC2 | 2022-11-02 | Bouldin Alfalfa-ADEX Heat Flux Plate Bouncer C | Removed | Lab none | Removed from about 2m south of the Bouncer tower at -10cm at the end of the 2022 advection experiment |
HP-127-1.0-1.3-71P PC1 | 2022-11-02 | Bouldin Alfalfa-ADEX Heat Flux Plate Bouncer B | Removed | Lab none | Removed from about 2m south of the Bouncer tower at -10cm at the end of the 2022 advection experiment |
HFP01 003072 | 2022-11-02 | Bouldin Alfalfa-ADEX Heat Flux Plate Bouncer A | Removed | Lab none | Removed from about 2m south of the Bouncer tower at -10cm at the end of the 2022 advection experiment |
LI-7500RS 75H-0128 | 2022-11-02 | Bouldin Alfalfa-ADEX Gas Analyzer Bouncer CO2/H2O | Removed | Lab none | Removed from top of the Bouncer tower at 3.96m alg at the end of the 2022 advection experiment at BA |
CR23X 2300 | 2022-11-02 | Bouldin Alfalfa-ADEX Data Logger Bouncer Met | Removed | Lab none | Removed from the Bouncer tower where it was collecting 30min averages from soil senors at the end of the 2022 advection experiment at BA |
LI-7500A 0073 | 2022-11-15 | Lab | Calibration | Lab | Last calibration: 2022-04-08 AGC=101.2 SDM= DAC1= DAC2= Ambient P=101.2kPa, Tair=23.2C, Tdew=1.69C Zero Air at 1850psi, AMP CA04119 524.77 at 1350 Tdewspan=20.00 CO2 zero 0.9163 -> .9165 Offset: .68ppm CO2 span 0.9983 -> 1.0085 Offset: -7.22ppm H2O zero 0.8246 -> .8296 Offset: 28.9ppm H2O span 1.0147 -> 1.0258 Offset: -.25ppm Notes: Used AMP LI-7550 sn AIU-1039 |
LI-7500A 0042 | 2022-11-15 | Lab | Calibration | Lab | Last calibration: 2022-04-19 AGC=99.1 SDM= DAC1= DAC2= Ambient P=101.2kPa, Tair=23.0C, Tdew=2.87C Zero Air at 1850psi, AMP CA04119 524.77 at 1350 Tdewspan=20.00 CO2 zero 0.9483 -> .9481 Offset: -.73ppm CO2 span 1.0087 -> 1.0127 Offset: -2.76ppm H2O zero 0.8444 -> .847 Offset: 16.3ppm H2O span 1.0191 -> 1.0292 Offset: -.22ppm Notes: Used AMP LI-7550 sn AIU-1039 |
LI-7500A 0041 | 2022-11-15 | Lab | Calibration | Lab | Last calibration: 2022-03-30 AGC=93.5 SDM= DAC1= DAC2= Ambient P=101.2kPa, Tair=22.9C, Tdew=.1C Zero Air at 1900psi, AMP CA04119 524.77 at 1350 Tdewspan=20.00 CO2 zero 0.9969 -> .9998 Offset: 9.35ppm CO2 span 1.0189 -> 1.0178 Offset: .23ppm H2O zero 0.8214 -> .8213 Offset: -.5ppm H2O span 1.0337 -> 1.0308 Offset: .07ppm Notes: Reset SS from 93.5 to 100 before calibrating. Used AMP LI-7550 sn AIU-1039. |
EXO fDOM 21K102591 | 2022-11-03 | Gilbert Tract | Calibration | Gilbert Tract | fDOM (QSU and RFU), Last cal: 10/31/2022 4:14:35 PM EXO sn 22C100485 EXO/Sensor Firmware: 1.0.86/3.0.0 UCSC sensors sn 591 and sn 592 were really off even though calibrated recently. Will accept calibration and just run them at the GT channel for a week to see if they drift. ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 QSU,0.51 QSU,0.01 QSU,0.00 QSU,16.967 °C Pt2,316.50 QSU,219.51 QSU,316.49 QSU,0.00 QSU,10.855 °C ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 RFU,0.12 RFU,0.00 RFU,0.00 RFU,15.878 °C Pt2,105.50 RFU,74.86 RFU,105.50 RFU,0.00 RFU,11.257 °C |
EXO fDOM 21K102592 | 2022-11-03 | Gilbert Tract | Calibration | Gilbert Tract | fDOM (QSU and RFU), Last cal: 10/31/2022 4:14:35 PM EXO sn 22C100485 EXO/Sensor Firmware: 1.0.86/3.0.0 UCSC sensors sn 591 and sn 592 were really off even though calibrated recently. Will accept calibration and just run them at the GT channel for a week to see if they drift. ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 QSU,1.87 QSU,0.00 QSU,0.00 QSU,16.796 °C Pt2,316.50 QSU,218.07 QSU,316.49 QSU,0.00 QSU,10.735 °C ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 RFU,0.20 RFU,0.00 RFU,0.00 RFU,15.801 °C Pt2,105.50 RFU,71.69 RFU,105.50 RFU,0.00 RFU,11.214 °C |
EXO fDOM 22C100780 | 2022-11-03 | Gilbert Tract | Calibration | Gilbert Tract | fDOM (QSU and RFU), Last cal: 10/24/2022 8:08:51 AM EXO sn 22C100485 EXO/Sensor Firmware: 1.0.86/3.0.0 UCSC sensors sn 591 and sn 592 were really off even though calibrated recently. Will accept calibration and just run them at the GT channel for a week to see if they drift. ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 QSU,-1.28 QSU,0.00 QSU,0.00 QSU,16.696 °C Pt2,316.50 QSU,315.87 QSU,316.54 QSU,0.00 QSU,10.444 °C ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 RFU,0.11 RFU,0.00 RFU,0.00 RFU,10.610 °C Pt2,105.50 RFU,92.32 RFU,105.51 RFU,0.00 RFU,11.275 °C |
EXO Turbidity 22C100485 | 2022-11-01 | Lab none | Calibration | Lab none | Turbidity (FNU), Last cal: Unknown EXO sn 22C100485 EXO/Sensor Firmware: 1.0.86/3.0.4 A factory reset was done first before this calibration was done in the lab. ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 FNU,0.24 FNU,0.00 FNU,0.00 FNU,20.384 °C Pt2,124.00 FNU,112.11 FNU,123.99 FNU,0.00 FNU,20.389 °C Pt3,1010.00 FNU,939.77 FNU,1009.99 FNU,0.00 FNU,20.565 °C |
EXO pH-ORP 22D102641 | 2022-09-23 | Lab none | Calibration | Lab none | ORP (mV), Last cal: Unknown EXO sn 22C103416 EXO/Sensor Firmware: 1.0.86/3.0.0 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,223.0 mV,220.6 mV,223.0 mV,0.0 mV,25.000 °C |
EXO pH-ORP 22D102641 | 2022-09-23 | Lab none | Calibration | Lab none | pH, Last cal: 7/12/2022 2:26:48 PM EXO sn 22C103416 EXO/Sensor Firmware: 1.0.86/3.0.0 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,7.00 pH,6.96 pH,7.00 pH,-4.1 mV,24.500 °C Pt2,10.01 pH,10.15 pH,10.01 pH,-181.2 mV,24.500 °C Pt3,4.01 pH,3.83 pH,4.02 pH,169.1 mV,24.500 °C |
EXO Turbidity 22E106072 | 2022-09-23 | Lab none | Calibration | Lab none | Turbidity (FNU), Last cal: Unknown EXO sn 22C103416 EXO/Sensor Firmware: 1.0.86/3.0.4 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 FNU,-0.19 FNU,0.00 FNU,0.00 FNU,24.000 °C Pt2,124.00 FNU,124.85 FNU,124.00 FNU,0.00 FNU,24.000 °C |
EXO fDOM 22E100649 | 2022-09-23 | Lab none | Calibration | Lab none | fDOM (QSU and RFU), Last cal: Unknown EXO sn 22C103416 EXO/Sensor Firmware: 1.0.86/3.0.0 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.00 QSU,-1.65 QSU,0.00 QSU,0.00 QSU,25.000 °C Pt2,308.00 QSU,301.85 QSU,307.99 QSU,0.00 QSU,15.000 °C ,Standard,PreCal,PostCal,RawValue,Temp Pt1,103.00 RFU,98.95 RFU,103.00 RFU,0.00 RFU,15.000 °C Pt2,0.00 RFU,-0.68 RFU,0.00 RFU,0.00 RFU,25.000 °C |
EXO2 22C103416 | 2022-09-23 | Lab none | Calibration | Lab none | Depth (m), Last cal: 6/27/2022 1:49:10 PM EXO sn 22C103416 EXO/Sensor Firmware: 1.0.86/3.1.3 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.000 m,-0.008 m,0.000 m,0.000 m,22.000 °C |
EXO pH-ORP 22D102641 | 2022-07-12 | Lab none | Calibration | Lab none | pH, Last cal: Unknown EXO sn 22C103416 EXO/Sensor Firmware: 1.0.86/3.0.0 Slope is lower than what it should be for a new sensor,Standard,PreCal,PostCal,RawValue,Temp Pt1,7.00 pH,7.02 pH,7.00 pH,-1.4 mV,20.000 °C Pt2,10.00 pH,9.89 pH,10.00 pH,-167.8 mV,20.000 °C Pt3,4.00 pH,4.39 pH,4.02 pH,151.1 mV,20.000 °C |
EXO2 22C103416 | 2022-06-27 | Lab none | Calibration | Lab none | Depth (m), Last cal: Unknown EXO sn 22C103416 EXO/Sensor Firmware: 1.0.86/3.1.3 ,Standard,PreCal,PostCal,RawValue,Temp Pt1,0.000 m,10.300 m,0.000 m,0.000 m,18.800 °C |
PAR-LITE 030382 | 2022-10-31 | Lab none | Tested | Lab none | This sensor was removed because there were spikes in its data on two different mux ports. I pulled apart the splice in the cable and didn't see any problems. From 2022-10-27 to 2022-10-31 I ran an inter-comparison out the lab window with sn 010164. The relationship was tight (r2=0.9999) and the slope was about 1.06. sn 030382 had been inter-compared on the Tonzi tower recently but in the reflected position so the range was not optimal. I think both sensors are functional and should be inter-compared with a good reference in the field to determine appropriate coefficients. |
PAR-LITE 010164 | 2022-10-31 | Lab none | Repair | Lab none | One of the wires was broken off right at the base of the sensor. I was able to solder on a lead and spliced a longer cable on and the sensor seemed to respond to light. From 2022-10-27 to 2022-10-31 I ran an inter-comparison out the lab window with sn 030382. The relationship was tight (r2=0.9999) and the slope was about 1.06. sn 030382 had been inter-compared on the Tonzi tower recently but in the reflected position so the range was not optimal. I think both sensors are functional and should be inter-compared with a good reference in the field to determine appropriate coefficients. |
CR23X 2143 | 2022-10-27 | Lab none | Disposed | Disposed none | I took this 23x apart and found a surface mount capacitor that had blown. I tried to repair the board trace and replaced the cap with one from a dead CR10x - no way to tell if the components were the same. After putting it back together it was still bricked. We will keep it for parts for a while. |
HYDROS21 H21G20001202 | 2022-10-19 | Lab none | Installed | Gilbert Tract Water Level | Installed at GT off of the boardwalk where it is co-located with the eosGP and Campbell conductivity sensor. When the staff gauge read 21cm, the bottom of the sensors was 19cm below the water. |
PA-II-SD 84f3ebd5ffb1 | 2022-10-18 | Bouldin Corn Air Quality Sensor | Removed | Lab none | Removed from BC tower and brought back to the lab for further inspection. Data over the summer seemed spiky. |
Mark II C013311 | 2022-10-18 | Lab none | Installed | Bouldin Alfalfa Weather Station C | Installed at BA (eastern position) to replace sn C003240, which was removed b/c of a broken T/RH sensor. |
Mark II C003240 | 2022-10-18 | Bouldin Alfalfa Weather Station C | Removed | Lab none | Removed from BA (eastern position) because its T/RH sensor is broken. Brought back to the lab to ship back to Arable. It was replaced by sn C013311. |
Mark II C013433 | 2022-10-18 | Lab none | Installed | Bouldin Alfalfa Weather Station B | Installed at BA (western position) to replace sn C006164, which was removed b/c of a broken T/RH sensor. |
Mark II C006164 | 2022-10-18 | Bouldin Alfalfa Weather Station B | Removed | Lab none | Removed from BA (western position) because its T/RH sensor is broken. Brought back to the lab to ship back to Arable. It was replaced by sn C013433. |
OS136-1-MVC NO1107536 | 2022-10-20 | Tonzi none | Removed | Lab none | This sensor has not been sending good data from a while. Seems a woodpecker destroyed its window. Removed and brought in for disposal |
SI-131 7385 | 2022-10-20 | Lab none | Installed | Tonzi Infrared Temperature Ref Panel A | Install at the Tonzi cal panel for calibration. But I damaged the data logger so not collecting data until next time. |
CR23X 2143 | 2022-10-20 | Lab none | Damaged | Lab none | I tried to install this at Tonzi to collect data from and IRT at the cal panel but fried it with 24V instead of 12V. It flashed "Hello" before something burned on the board. |
PAR-LITE 990025 | 2022-10-20 | Lab none | Installed | Tonzi PAR Sensor Reflected | Installed at Tonzi to replace sn 030382 that has been spiky. |
PAR-LITE 030382 | 2022-10-20 | Tonzi PAR Sensor Reflected | Removed | Lab none | Removed from Tonzi due to random spikiness. Replaced by sn 990025. |
SI-131 7385 | 2022-10-19 | Experiment none | Removed | Lab none | I removed this sensor from the bike kit to calibrate it against the cal panel at Tonzi. |
PAR-LITE 990025 | 2022-10-19 | Experiment none | Removed | Lab none | I removed this sensor from the bike kit to use at Tonzi. I also spliced about 3m of extra length on to its cable. |
CR23X 2143 | 2022-10-19 | Experiment none | Removed | Lab none | I removed this data logger from the bike kit and prepared it to install at Tonzi to do a calibration on the IRT. |
CTD-10 1782303006 | 2022-10-18 | Factory none | Disposed | Disposed none | Apparently the sensor was unrepairable and a new sensor was sent to us a replacement since this sensor was still under warranty. No info on the defect. The new sensor is a new model Hydros21 sn H21G200001202. |
CTD-10 1782303006 | 2022-09-20 | Lab none | Shipped | Factory none | After testing in the lab reproduce the errors seen in the field an RMA-008376 was received and this sensors was shipped to METER for repair - it should be under warranty. |
SD130BN 0030F4D214A4 | 2022-10-07 | Factory none | Returned | Lab none | Returned from Stardot after inspection determined that a power surge on the ethernet port cooked the innards. Will use the lens on a new body |
GMP222 K1310102 | 2022-10-05 | Tonzi-SoilCO2 CO2 Probe Open Profile -0.16 | Removed | Lab none | The soil co2 gradient system has not been in good condition for a long time. Finally getting around to removing the equipment from the field: Open -16cm |
GMP221 K1340006 | 2022-10-05 | Tonzi-SoilCO2 CO2 Probe Open Profile -0.08 | Removed | Lab none | The soil co2 gradient system has not been in good condition for a long time. Finally getting around to removing the equipment from the field: Open -8cm |